Larkspur Inc. issues 500 shares of $10 par value common stock and 100 shares of $100 par value preferred stock for a lump sum of $104,000.

(a) Prepare the journal entry for the issuance when the market price of the common shares is $164 each and market price of the preferred is $205 each.
(b) Prepare the journal entry for the issuance when only the market price of the common stock is known and it is $178 per share.
Answer #1
Answer

Requirement ‘a’

Accounts title Debit Credit
Cash $     104,000
Common Stock $       5,000
Paid in Capital in excess of Par - Common Stock $     78,200
Preferred Stock $     10,000
Paid in Capital in excess of Par - Preferred Stock $     10,800
(To record Stock issued for cash)
--Working
No. of shares Market Price Total Percent of total
Common Stock 500 $          164 $82,000 80.00%
Preferred Stock 100 $          205 $20,500 20.00%
Total fair Market Value $102,500 100.00%
Common Stock Preferred Stock
A Total Lumpsum Issue Price $104,000 $104,000
B Allocation % 80% 20%
C = A x B Total $83,200 $20,800
D No. of shares 500 100
E Par Value $10 $100
F = D x E Capital Stock credited by $5,000 $10,000
G = C - F Paid in Capital credited by $78,200 $10,800
Requirement ‘b’
Accounts title Debit Credit
Cash $         104,000
Common Stock $      5,000
Paid in Capital in excess of Par - Common Stock $    84,000
Preferred Stock $    10,000
Paid in Capital in excess of Par - Preferred Stock $      5,000
( To Record Stock issued for cash)
--Working
A Total Lumpsum $104,000
Common Stock:
B Market price of Common Stock $178
C No. of common stock shares issued 500
D = B x C Amount allocated to Common Stock $89,000
E = C x $ 10 par Common Stock account [$10 par] $5,000
F = D - E Paid in Capital credited by $84,000
Preferred Stock:
G = A - D Remaining amount for Preferred Stock $15,000
H No. of preferred Stock issued 100
I = H x $ 100 par Preferred Stock account [at $ 100 par] $10,000
J = G - I Paid in Capital credited by $5,000